### v2

@@ 4 Conclusions
  • a data-driven basis $\{\phi_{j}\}_{j=1}^{N}$ whose span $V=\operatorname{span}\{\phi_{j}\}_{j=1}^{N}$ is approximately invariant under $\mathcal{L}$.
SABON cleanly achieves this by combining an encoding–reconstruction structure with a linear latent map $\mathcal{G}$ and a composite loss that balances operator approximation accuracy and sparsity. We proved a universal approximation theorem in the spirit of Hua [16], specialised to a single neural basis that targets transfer and Koopman operators.
Two canonical examples were used to highlight key features of SABON:
@@
    Circle rotation. The SABON-learned basis accurately reconstructed the analytic eigenpairs and recovered the rotation angle $\alpha$ directly from the leading eigenvalues, while the associated eigenfunctions coincided with the theoretical ${e^{\,ik\theta}}$ modes. Thus, the learned basis accurately captured the main dynamical properties.
@@
    Perturbed Arnold’s cat map. SABON’s learned basis outperformed a traditional Fourier basis for a nonlinear uniformly hyperbolic map by adapting to the geometry inherent in the dynamics. This is particularly surprising given the well-known theoretical approximation strengths of Fourier modes. The network discovered basis functions largely aligned with the contracting and expanding directions, leading to more accurate estimates of the physical invariant measure of the dynamics and a highly invariant approximation subspace.
@@ Acknowledgments
The research of GF is supported by an Australian Research Council (ARC) Laureate Fellowship FL230100088. The research of KK is supported by a UNSW University International Postgraduate Award, with additional funding from UNSW’s School of Mathematics and Statistics and the ARC Laureate Fellowship. Computational resources were generously provided by Katana infrastructure at UNSW [30], which significantly facilitated this research.

### v3

@@ 4 Conclusions
  • a data-driven basis $\{\phi_{j}\}_{j=1}^{N}$ whose span $V=\operatorname{span}\{\phi_{j}\}_{j=1}^{N}$ is adapted to the underlying dynamics and provides an efficient finite-dimensional representation for $\hat{\mathcal{L}}$ and other approximations of $\mathcal{L}$.
SABON cleanly achieves this by combining an encoding–reconstruction structure with a linear latent map $\mathcal{G}$ and a composite loss that balances operator approximation accuracy and sparsity. We proved a universal approximation theorem in the spirit of Hua [HUA202321], specialised to a single neural basis that targets transfer and Koopman operators.
Three canonical examples were used to highlight key features of SABON:
@@
    Circle rotation. The SABON-learned basis accurately reconstructed the analytic eigenpairs and recovered the rotation angles $\alpha, kk\in\mathbb{Z}$, directly from the leading eigenvalues, while the associated eigenfunctions coincided with the theoretical $e^{\,ik\theta}$} modes. Thus, the learned basis accurately captured the main dynamical properties. In this setting, where the training data lies in a finite-dimensional invariant subspace, SABON recovers an invariant basis.
@@
    Weakly nonlinear perturbed cat map. For this uniformly hyperbolic map with approximately linear stable and unstable foliations, the learned basis adapted to the geometry of the dynamics and outperformed a Fourier basis of the same size when approximating the SRB measure.
  3. 3.
    Strongly nonlinear conjugated cat map. For a more geometrically demanding Anosov map with curved foliations and significant deviation from area preservation, SABON continued to produce basis functions adapted to the nonlinear geometry and outperformed Fourier bases. This is particularly surprising given the well-known theoretical approximation strengths of Fourier modes. The network discovered basis functions largely aligned with the contracting and expanding directions, leading to more accurate estimates of the physical invariant measure of the dynamics.
@@ Acknowledgments
The authors thank Irteza Chaudhry for helpful discussions and for pointing out an error in an earlier version. The research of GF is supported by an Australian Research Council (ARC) Laureate Fellowship FL230100088. The research of KK is supported by a UNSW University International Postgraduate Award, with additional funding from UNSW’s School of Mathematics and Statistics and the ARC Laureate Fellowship. Computational resources were generously provided by Katana infrastructure at UNSW [katana], which significantly facilitated this research.